### v1

@@ 7.3 Active learning using BMLP
Input: Hypotheses $H$, background knowledge $BK$, experiment instances $T$, experimental cost budget $C$ and oracle $O$.
@@
2:  while $cost(T_{selected})\leq C$ and $|V_{j}|>1$ do
@@ 8 Experiments
In this work, we define the experimental resource cost function as the cost of the reagents (optional nutrient substances) used in selected experiments. We also assume when added to growth media, reagents are used in the same volume. In addition, we do not account for plastic consumables, equipment deprecation or time to carry out each experiment, which we consider fixed. We also do not account for costs incurred due to different experiment types such as different measurement timescales that prevent experiments from being performed concurrently. To maintain consistent scaling for comparisons, the total experimental reagent cost is the amount of unit cost required by selected experiments. We consider the unit cost of the reagents as the cost of the cheapest reagent. We calculate the total experimental reagent cost as the sum of the reagent costs [^3] We provide the costs of reagents in the Appendix. in selected experiments divided by the unit cost.
@@
Using only Prolog to simulate metabolic networks requires grounding viable reaction pathways based on SLD-resolutions. We aim to demonstrate that a general-purpose Prolog interpreter can be augmented with our BMLP approach for faster phenotypic simulations given logically encoded metabolic network models. $BMLP_{active}$ extends the logical representation of a metabolic network in the Robot Scientist King04:RobotScientist ; ase_progol by additionally applying boolean matrix computation via BMLP. Therefore, we consider the Robot Scientist’s logical encoding as a benchmark. Since the Robot Scientist was implemented using Progol progol , we adapt its metabolic network encoding to SWI-Prolog. To answer Q1, we compare the simulation runtime of $BMLP_{active}$, which uses SWI-Prolog and BMLP-IE, against the Robot Scientist’s encoding in SWI-Prolog.
@@ 8.1 Experiment 1: BMLP-IE simulations and predictions
Table 1: Simulation run-times in CPU time. BMLP-IE leads to a 170 times improvement in simulation time efficiency. Multi-threading BMLP-IE enhances runtime efficiency 600 times compared to base SWI-Prolog.
@@ 8.2 Experiment 2: active learning sample complexity and experiment cost
Materials. Up to 3 of 16 common carbon sources were used as additional nutrients to the base growth medium. Each of these 16 carbon sources is an optional nutrient in the experimental data for validating iML1515 . This set contains preferred carbon sources martinez-gomez_new_2012 ; zampieri_regulatory_2019 for E. coli such as glucose and glycerol. We first simulated iML1515 with up to 3 optional carbon sources, generating $\sum_{0}^{i=3}\binom{16}{i}=697$ synthetic data. This allows us to go beyond the 16 experimental data for each gene knockout in iML1515 . We observed a single-knockout phenotypic effect from 213 genes. We treated these synthetic data as ground truths where 213 genes have only positive examples. A subset of 3 genes has both positive and negative examples [^4] The distribution of training examples for this experiment is described in the Appendix.. We focused on these 3 genes. We removed gene functions associated with these 3 genes from the background knowledge. Thus, the resulting incomplete model has a smaller set of $codes/2$ clauses.
Methods. We focused on one removed gene function at a time to recover it. Each hypothesis except the empty hypothesis associates a gene locus identifier, whose function we tried to recover, with one of the 1294 non-maintenance reactions in iML1515. Therefore, the hypothesis space had 1295 hypotheses including an empty hypothesis and gene function hypotheses ($codes/2$ clauses). The instance space contained 697 experiments with optional growth nutrients up to 3 carbon sources.
@@
$BMLP_{active}$ actively selected experiments from the instance space to reduce the hypothesis space. The random selection method randomly sampled instances from the instance space. After selecting experiments, both methods returned the hypothesis with the single highest compression, otherwise, a hypothesis was selected from those sharing the highest compression randomly. The final hypotheses obtained by either method would output a recovered model. We compared $BMLP_{active}$ with the random experiment selection strategy with increasing experimental reagent cost budget {1, 10, 50, 100, 500, 1000, 5000, 10000}. We tested the predictive accuracy of the recovered models against all 697 synthetic data. The cost of selected experiments and predictive accuracy of the recovered model were recorded for $BMLP_{active}$ and random selection. We show this comparison in Fig 5.
@@
We also compared $BMLP_{active}$ to random experiment selection with upper bounds on the number of selected experiments from {0, 1, 2, 3, 4, 5, 10, 15, 20, 25, 30}. Each selection method was repeated 10 times given a reagent cost budget or an upper bound on the number of experiments. In addition, we show this comparison in Fig 6. We computed the phenotypic predictive accuracy of the recovered models against the iML1515 background knowledge on experiments with up to 3 carbon sources. For experiment number limits excluding 0, we calculated the mean error of learned hypotheses and divided it by the number of experiments selected to obtain the ratio. In Fig 7, we demonstrate how these ratios change to support Theorem 6.1.
@@
Results and discussion. Fig 5 shows that $BMLP_{active}$ spends 10% of the cost used by random experiment sampling when converging to the correct hypothesis. The y-axis is the total experimental reagent cost in $log_{10}$ calculated from selected experiments. And 100% predictive accuracy indicates successful recovery of the deleted gene functions. The correct gene functions were recovered by $BMLP_{active}$ and random experiment selection with $10^{1.99}$ and $10^{2.99}$ experimental reagent costs respectively. However, $BMLP_{active}$ only used $\frac{1}{10}$ the cost of random experiment selection (grey dotted lines). The total experimental reagent cost is computed by dividing the unit cost which is the cost of the cheapest optional nutrient at £0.038. Therefore, conducting experiments selected by $BMLP_{active}$ and random sampling would cost $10^{1.99}\times 0.038\approx\pounds 3.8$ and $10^{2.99}\times 0.038\approx\pounds 38$ to learn a single gene function, given the same scale of hypothesis space and experimental design space. Consideration of differential resource factors in alternative user-defined experimental cost functions could significantly increase the total experimental resource cost.
@@
We then examined the advantage of $BMLP_{active}$ over random experiment selection on the number of experiments needed to recover gene functions. The number of experiments is a significant determinant of experimental cost. Fig 6 shows that $BMLP_{active}$ reduces the number of instances needed for labelling to learn accurate gene functions. While the random experiment selection method requires 25 experiments to recover the deleted gene functions, $BMLP_{active}$ only needed 3 experiments. Fig 5 and Fig 6 answer the question Q3 since $BMLP_{active}$ successfully reduced the experimental reagent cost and experiment cost compared to random experiment selection. Fig 5 and Fig 6 show $BMLP_{active}$ can simultaneously optimise a user-defined experimental resource cost function and reduce the total number of experiments to learn gene functions accurately. This means that $BMLP_{active}$ is highly flexible and could be tailored for specific experimental objectives in a discovery process. Later, in Experiment 3, we revisit the question Q3 for recovering a known gene-isoenzyme association which presents a bigger empirical challenge due to the need to explore a larger experimental design space.
@@
Fig 7 demonstrates the mathematical relationships between predictive error and the number of experiments required by our active learner $BMLP_{active}$ and random experiment selection to converge to the correct hypothesis. The ratio $\frac{\epsilon}{s_{active}}$ is greater than $\frac{\epsilon+\phi}{s_{passive}}$ for $\epsilon>0$ and $\phi=0.5\%$. There are fewer points on the active learning ratio curve since $BMLP_{active}$ converges to 100% accuracy using fewer training examples. The two curves intersect at $\epsilon=0$ since both converge to 100% accuracy. Fig 7 answers the question Q4 and shows that Theorem 6.1 correctly describes the relationship between the sample complexity of our active learner and random experiment selection. Theorem 6.1 supports that active learning can reduce the requirement for training examples to achieve a certain learning performance. In the context of scientific discovery, the experimental resource cost to arrive at a finding can be reduced by active learning since each actively selected experiment can have a higher return of information.
@@ 8.3 Experiment 3: learning digenic functions
Materials. We focused on re-discovering a target isoenzyme function associated with the gene $tyrB$ in the Tryptophan biosynthesis pathways. We aim to demonstrate gene-isoenzyme function learning by recovering this gene function. Associated functions of $tyrB$ and $aspC$ can be classified as isoenzyme functions since they are responsible for producing an important amino acid L-phenylalanine. In our logical representation of iML1515, we use reaction and $codes/2$ unit ground literals to describe an isoenzyme. We selected 33 genes associated with this essential pathway. Aromatic amino acids are fundamental for protein synthesis but also serve as precursors for vital secondary metabolites and high-value chemicals choi_violacein_2015 . We consider six aromatic amino acids as optional nutrients for recovering the deleted $tyrB$ isoenzyme function.
@@
Methods. We removed tyrB’s gene function and metabolic reaction associated with this isoenzyme from $BMLP_{active}$. The experiment instance space contained double gene-knockout and single gene-knockout experiments of 33 genes. There were $\binom{33}{2}$ gene pair combinations for double-gene knockout experiments and 33 single-gene knockout experiments. We considered 7 growth medium conditions. Therefore, the experiment instance space was $(\binom{33}{2}+33)\times 7=3927$. We used the iML1515 background knowledge to generate synthetic data from 3696 double-knockout experiments. In addition, we used 231 single-knockout empirical data in iML1515 .
@@
In the metabolic network, an isoenzyme function hypothesis includes a node represented by a new enzyme identifier (a $enzyme/6$ clause)and a new link between a gene locus identifier and this new enzyme identifier (a $codes/2$ clause). The hypothesis space was defined by the associations of 33 genes (from the focused library) with 33 enzymes and reactions. There are $33\times 32+1=1057$ hypotheses including potential gene-isoenzyme functions and an empty hypothesis. We increased linearly and exponentially the number of experiments $N$ that could be selected according to {1, 2, 4, 5, 8, 10, 15, 16, 20, 25, 30, 32, 64, 128, 256}. $BMLP_{active}$ selected $N$ experiments from this instance space to actively learn from the hypothesis space. The random selection strategy randomly sampled $N$ instances from the instance space. Again, both methods output the hypothesis with the highest compression according to the sampled examples. A hypothesis was randomly selected when multiple competing hypotheses shared the highest compression. We recovered the models by adding the final hypotheses. We recorded the number of experiments selected and evaluated if the recovered models contained the target isoenzyme function from the iML1515 background knowledge. Each experiment selection method was repeated 10 times and we computed the frequency of successful isoenzyme recovery across these 10 repeats. We show this comparison in Fig 8.
@@
This result also shows that $BMLP_{active}$ significantly reduced the number of experiments needed compared to random experiment selection, which answers the question Q3. $BMLP_{active}$ provides higher information gain from each experiment and can guarantee recovery with as few as 20 experiments. While random sampling could prune most candidate hypotheses, it failed to eliminate competitive hypotheses further. In contrast to $BMLP_{active}$, random sampling could not recover this isoenzyme function with 64 experiments and could only do this occasionally with more than 10 times the number of experiments required by $BMLP_{active}$. Although we used a reduced set of 30 genes in this demonstration, the combinatorial space was sufficiently large that random experiment selection became non-viable as an experimentation strategy in a discovery process. The remarkable increase in efficiency with $BMLP_{active}$ with this task demonstrates it has potential even as the experimental design space grows.
@@ 9 Conclusion and future work
We present a new logic programming system $BMLP_{active}$ that can simulate and learn gene functions in a state-of-the-art GEM. This GEM, iML1515, describes 1515 genes and 2719 metabolic reactions in E. coli. To our knowledge, $BMLP_{active}$ is the first logic programming system able to operate on GEMs. We implement $BMLP_{active}$ via our proposed BMLP approach by exploiting boolean matrices to improve the simulation runtime of a GEM by 170 times. Our empirical results show that $BMLP_{active}$ can learn gene functions while reducing the resource cost of experimentation by 90%. In addition, we deploy $BMLP_{active}$ to actively learn gene-isoenzyme functions which are an error source in GEM. As a demonstration, we empirically show that $BMLP_{active}$ can learn an important isoenzyme function in the aromatic amino acid biosynthesis pathway with as little as 20 training examples. In comparison, random experiment selection cannot this isoenzyme function consistently even with more than 250 training data. $BMLP_{active}$ provides a substantial advantage for isoenzyme function discovery over random experiment selection due to the sample complexity reduction. This finding suggests that $BMLP_{active}$ could be used to offset the growth in experimental design spaces associated with genome-scale function learning.
@@
Designing experiments to explain some hypothesised phenomenon requires deliberate thinking by scientists. A helpful scientific assistant could teach scientists to create experiment plans with optimal outcomes. This aspect of human-AI collaboration has been studied in an emerging research area on Ultra-Strong Machine Learning (USML) michie1988 . A USML not only learns a concept but also teaches this concept to humans such that human comprehension is improved. Prior research has shown instances of USML in the context of learning recursive concepts US2018 and a two-player game Ai2021 . A recent study sequential_teaching demonstrated that an ILP system can play the role of a “teacher” and learning from machine explanations of its output can inspire human discovery. To date, there is no demonstration of autonomous discovery systems that satisfy the USML criterion. The ultimate goal of autonomous discovery systems is to acquire new scientific understanding and explain insights to humans krenn_scientific_2022 . The present work can be extended to examine the effect of showing actively selected experiments on human experiment selection and their comprehension of complex biological concepts.
@@
Datasets used in empirical studies do not contain any personal information or other sensitive content. There are no ethical issues.

### v3

@@ 7.3 Active learning using BMLP
Input: Hypotheses $H$, background knowledge $BK$, experiment instances $T$, experimental cost budget $C$, total number of experiment $N$ and oracle $O$.
@@
2:  while $cost(T_{selected})\leq C$ and $|V_{j}|>1$ and $|T_{selected}|<N$ do
@@ 8 Experiments
In this work, we define the experimental cost function as the cost of the reagents which are optional nutrient substances used in selected experiments. We also assume when added to growth media, reagents are used in the same volume. In addition, we do not explicitly account for plastic consumables, equipment deprecation or specify the time to carry out each experiment, which we consider fixed and are affected mainly by the number of experiments. We also do not account for costs incurred due to different experiment types such as different measurement timescales that prevent experiments from being performed concurrently. To maintain consistent scaling for comparisons, the total experimental reagent cost is the amount of unit cost required by selected experiments. We consider the unit cost of the reagents as the cost of the cheapest reagent. We calculate the total experimental reagent cost as the sum of the reagent costs [^4] We provide the costs of reagents in Table 2 and 3 of Appendix A. in selected experiments divided by the unit cost.
@@
Using only Prolog to simulate metabolic networks requires grounding viable reaction pathways based on SLD resolutions. We aim to demonstrate that a general-purpose Prolog interpreter can be augmented with our BMLP approach for faster phenotypic simulations given logically encoded metabolic network models. $BMLP_{active}$ extends the logical representation of a metabolic network in the Robot Scientist King04:RobotScientist ; ase_progol by additionally applying boolean matrix computation via BMLP. Therefore, we consider the Robot Scientist’s logical encoding as a benchmark. Since the Robot Scientist was implemented using Progol progol , we adapt its metabolic network encoding to SWI-Prolog. To answer Q1, we compare the simulation runtime of $BMLP_{active}$, which uses SWI-Prolog and BMLP-IE, against the Robot Scientist’s encoding in SWI-Prolog.
@@ 8.1 Experiment 1: BMLP-IE simulations and predictions
Table 1: Simulation runtimes in CPU time. BMLP-IE leads to a 170 times improvement in simulation time efficiency. Multi-threading BMLP-IE enhances runtime efficiency 600 times compared to base SWI-Prolog.
@@ 8.2 Experiment 2: active learning sample complexity and experiment cost
Materials. Up to 3 of 16 common carbon sources were used as additional nutrients to the base growth medium. Each of these 16 carbon sources is an optional nutrient in the experimental data for validating iML1515 . This set contains preferred carbon sources martinez-gomez_new_2012 ; zampieri_regulatory_2019 for E. coli such as glucose and glycerol. We first simulated iML1515 with up to 3 optional carbon sources, generating $\sum_{0}^{i=3}\binom{16}{i}=697$ synthetic data. This allows us to go beyond the 16 experimental data for each gene knockout in iML1515 . We observed a single-knockout phenotypic effect from 213 genes. We treated these synthetic data as ground truths where 213 genes have only positive examples. We only focused on a subset of 2 genes because they have both positive and negative examples [^5] The distribution of training examples for this experiment is described in Table 4 of Appendix B.. Then, we removed gene functions associated with these 2 genes from the background knowledge. Thus, the resulting incomplete model has a smaller set of $codes/2$ clauses.
Methods. We focused on one removed gene function at a time to recover it. Each hypothesis except the empty hypothesis associates a gene locus identifier, whose function we tried to recover, with an $enzyme/6$ clause. We concentrate on 1294 $enzyme/6$ clauses that represent the non-maintenance reactions in iML1515. Therefore, the hypothesis space had 1295 $codes/2$ hypotheses, including an empty hypothesis. The instance space contained 697 experiments with optional growth nutrients up to 3 carbon sources.
@@
$BMLP_{active}$ actively selected experiments from the instance space to reduce the hypothesis space. The random selection method randomly sampled instances from the instance space. First, we constrained the selection of experiments by an experimental reagent cost budget $C$. After experiment selection, both methods returned the hypothesis with the single highest compression, otherwise, a hypothesis was selected from those sharing the highest compression randomly. The final hypotheses obtained by either method would output a recovered model. We compared $BMLP_{active}$ with the random experiment selection strategy with increasing experimental reagent cost budget $C\in\{1,10,50,100,500,1000,5000,10000\}$. We tested the predictive accuracy of the recovered models against all 697 synthetic data. The cost of selected experiments and predictive accuracy of the recovered model were recorded for $BMLP_{active}$ and random selection. Each selection method was repeated 10 times for each reagent cost budget $C$. We show this comparison in Fig 5.
@@
In addition, we constrained the number of experiments $N$ that can be selected by increasing it linearly according to $\{0,1,2,3,4,5,10,15,20,25,30\}$. We applied $BMLP_{active}$ and random experiment selection with compression to decide the final hypothesis for recovering models. Each selection method was repeated 10 times for every number of selected experiments $N$. Then, we computed the phenotypic predictive accuracy of the recovered models against all 697 synthetic data. We show this comparison in Fig 6. For experiment number limits excluding 0, we calculated the mean error of learned hypotheses and divided it by the number of experiments selected to obtain the ratio. In Fig 7, we demonstrate how these ratios change to support Theorem 6.1.
@@
Results and discussion. Fig 5 shows that $BMLP_{active}$ spends 10% of the cost used by random experiment sampling when converging to the correct hypothesis. The y-axis is the total experimental reagent cost in $log_{10}$ calculated from selected experiments. And 100% predictive accuracy indicates successful recovery of the deleted gene functions. The correct gene functions were recovered by $BMLP_{active}$ and random experiment selection with $10^{1.99}$ and $10^{2.99}$ experimental reagent costs respectively. The total experimental reagent cost is computed by dividing the unit cost which is the cost of the cheapest optional nutrient (£$\pounds 0.038/gram$ ). Given the current hypothesis space and experimental design space and assuming each reagent was used one gram, experiments selected by $BMLP_{active}$ and random sampling would cost $10^{1.99}\times 0.038\approx\pounds 3.8$ and $10^{2.99}\times 0.038\approx\pounds 38$ to learn a single gene function. Consideration of differential resource factors in the user-defined experimental cost function could significantly increase the total experimental cost.
@@
We then examined the advantage of $BMLP_{active}$ over random experiment selection on the number of experiments needed to recover gene functions. The number of experiments is a significant determinant of experimental cost. Fig 6 shows that $BMLP_{active}$ reduces the number of instances needed for labelling to learn accurate gene functions. While random experiment selection requires 25 experiments to recover the deleted gene functions, $BMLP_{active}$ only needed 3 experiments. Fig 5 and Fig 6 answer the question Q3 since $BMLP_{active}$ successfully reduced the experimental reagent cost and experiment cost compared to random experiment selection. Fig 5 and Fig 6 show $BMLP_{active}$ can simultaneously optimise a user-defined experimental cost function and reduce the total number of experiments to learn gene functions accurately. This means that $BMLP_{active}$ is highly flexible and could be tailored for specific experimental objectives in a discovery process. Later, in Experiment 3, we revisit the question Q3 for recovering a known gene-isoenzyme association which presents a bigger empirical challenge due to the need to explore a larger experimental design space.
@@
Fig 7 demonstrates the mathematical relationships between predictive error and the number of experiments required by $BMLP_{active}$ and random experiment selection to converge to the correct hypothesis. For this gene function recovery task, $BMLP_{active}$ selected experiments with minimal reduction ratio $\phi\geq 0.5\%$. The ratio $\frac{\epsilon}{s_{active}}$ is greater than $\frac{\epsilon+\phi}{s_{passive}}$ for error $\epsilon>0$. There are fewer points on the active learning ratio curve since $BMLP_{active}$ converges to 100% accuracy using fewer training examples. The two curves intersect at $\epsilon=0$ since both converged to 100% accuracy. Fig 7 answers the question Q4 and supports Theorem 6.1. Theorem 6.1 shows that active learning ($BMLP_{active}$) can reduce the requirement for training examples compared with passive learning (random experiment selection) to achieve a certain learning performance. In the context of scientific discovery, the experimental cost to arrive at a finding can be reduced by active learning since each actively selected experiment can have a higher return of information.
@@ 8.3 Experiment 3: learning digenic functions
Materials. We focused on re-discovering a target isoenzyme function associated with the gene $tyrB$ in the Tryptophan biosynthesis pathways. We aim to demonstrate gene-isoenzyme function learning by recovering this gene function. Associated functions of $tyrB$ and $aspC$ can be classified as isoenzyme functions since they are responsible for producing an important amino acid L-phenylalanine. In our logical representation of iML1515, we use reaction and $codes/2$ unit ground literals to describe an isoenzyme. We selected 33 genes associated with this essential pathway. In this focus gene set, 27 genes are associated with a single function and 6 genes are related to two functions. We consider medium conditions with 7 optional nutrients including aromatic amino acids for recovering the deleted $tyrB$ isoenzyme function. Aromatic amino acids are fundamental for protein synthesis but also serve as precursors for vital secondary metabolites and high-value chemicals choi_violacein_2015 .
@@
Methods. We removed tyrB’s gene function and metabolic reaction associated with this isoenzyme from $BMLP_{active}$. The experiment instance space contained double gene-knockout and single gene-knockout experiments of 33 genes. There were $\binom{33}{2}$ gene pair combinations for double-gene knockout experiments and 33 single-gene knockout experiments. We considered growth medium conditions with the 7 optional nutrients. Therefore, the experiment instance space was $(\binom{33}{2}+33)\times 7=3927$ and 3696 experiment instance labels were synthetically generated from double-knockout experiments using the full iML1515 background knowledge. The rest 231 single-knockout empirical data were from validation data in iML1515 .
@@
In the metabolic network, an isoenzyme function hypothesis contains a node represented by a new enzyme identifier (a $enzyme/6$ clause) and a new link between a gene locus identifier and this new enzyme identifier (a $codes/2$ clause). The hypothesis space contained associations of 33 genes (from the focused library) with 33 $enzymes/6$ clauses. There were $(27\times 32+6\times 31)=1050$ potential hypotheses for 27 single-function genes and 6 double-function genes. We also included the tyrB function hypothesis and an empty hypothesis, which gave 1052 hypotheses in total. We increased linearly and exponentially the number of experiments $N$ that could be selected according to {1, 2, 4, 5, 8, 10, 15, 16, 20, 25, 30, 32, 64, 128, 256}. $BMLP_{active}$ selected $N$ experiments from this instance space to actively learn from the hypothesis space. The random selection strategy randomly sampled $N$ instances from the instance space. Again, both methods output the hypothesis with the highest compression according to the sampled examples. A hypothesis was randomly selected when multiple competing hypotheses shared the highest compression. We recovered the models by adding the final hypotheses. We recorded the number of experiments selected and evaluated if the recovered models contained the target isoenzyme function from the iML1515 background knowledge. Each experiment selection method was repeated 10 times and we computed the frequency of successful isoenzyme recovery across these 10 repeats. We show this comparison in Fig 8.
@@
This result also shows that $BMLP_{active}$ significantly reduced the number of experiments needed compared to random experiment selection, which answers the question Q3. $BMLP_{active}$ provides higher information gain from each experiment and can guarantee recovery with as few as 20 experiments. While random sampling could prune most candidate hypotheses, it failed to eliminate competitive hypotheses further. In contrast to $BMLP_{active}$, random sampling could not recover this isoenzyme function with 64 experiments and could only do this occasionally with more than 10 times the number of experiments required by $BMLP_{active}$. Although we used a reduced set of genes in this demonstration, the combinatorial space was sufficiently large that random experiment selection became non-viable as an experimentation strategy in a discovery process. The remarkable increase in efficiency with $BMLP_{active}$ with this task demonstrates it has potential even as the experimental design space grows.
@@ 9 Conclusion and future work
We present a new logic programming system $BMLP_{active}$ that can simulate and learn gene functions in a state-of-the-art GEM. This GEM, iML1515, describes 1515 genes and 2719 metabolic reactions in E. coli. To our knowledge, $BMLP_{active}$ is the first logic programming system able to operate on GEMs. We implement $BMLP_{active}$ via our proposed BMLP approach by exploiting boolean matrices to improve the simulation runtime of a GEM by 170 times. Our empirical results show that $BMLP_{active}$ can learn gene functions while reducing the cost of optional nutrient substance in experiments by 90%. In addition, we deploy $BMLP_{active}$ to actively learn gene-isoenzyme functions which are an error source in GEM. As a demonstration, we empirically show that $BMLP_{active}$ can learn an important isoenzyme function in the aromatic amino acid biosynthesis pathway with as little as 20 training examples. In comparison, random experiment selection cannot learn this isoenzyme function consistently even with more than 250 training data. $BMLP_{active}$ provides a substantial advantage for isoenzyme function discovery over random experiment selection due to the sample complexity reduction. This finding suggests that $BMLP_{active}$ could be used to offset the growth in experimental design spaces associated with genome-scale function learning.
@@
Designing experiments to explain some hypothesised phenomenon requires deliberate thinking by scientists. A helpful scientific assistant could teach scientists to create experiment plans with optimal outcomes. This aspect of human-AI collaboration has been studied in an emerging research area on Ultra-Strong Machine Learning (USML) michie1988 . A USML not only learns a concept but also teaches this concept to humans such that human comprehension is improved. Prior research has shown instances of USML in the context of learning recursive concepts US2018 and a two-player game Ai2021 . A recent study sequential_teaching demonstrated that an ILP system can play the role of a “teacher” and human learning from machine explanations of its output can sometimes inspire human discovery. To date, there is no demonstration of autonomous discovery systems that satisfy the USML criterion. The ultimate goal of autonomous discovery systems is to acquire new scientific understanding and explain insights to humans krenn_scientific_2022 . The present work can be extended to examine the effect of machine-selected experiments and machine-learned output on human experiment selection and human comprehension of complex biological concepts.
@@
Datasets used do not contain any personal information or other sensitive content. There are no ethical issues.
Data Availability
Data used and generated in all experimental sections are accessible on
https://github.com/lai1997/BMLP_active_public.